# Sparse Polyak: an adaptive step size rule for high-dimensional M-estimation

### v1

@@ Abstract
We propose and study Sparse Polyak, a variant of Polyak’s adaptive step size, designed to solve high-dimensional statistical estimation problems where the problem dimension is allowed to grow much faster than the sample size. In such settings, the standard Polyak step size performs poorly, requiring an increasing number of iterations to achieve optimal statistical precision-even when, the problem remains well conditioned and/or the achievable precision itself does not degrade with problem size. We trace this limitation to a mismatch in how smoothness is measured: in high dimensions, it is no longer effective to estimate the Lipschitz smoothness constant. Instead, it is more appropriate to estimate the smoothness restricted to specific directions relevant to the problem (restricted Lipschitz smoothnes constant). Sparse Polyak overcomes this issue by modifying the step size to estimate the restricted Lipschitz smoothness constant. We support our approach with both theoretical analysis and numerical experiments, demonstrating its improved performance.
@@ 3 Main Result
This theorem focuses on the distance of the iterates to $\theta^{*}$. The quantity $T$ can be interpreted as the number of iterations required to reach the desired accuracy when applying Algorithm 1 with $\widehat{f}=f^{*}$. The additional term $\mathcal{O}!\left(\frac{f(\theta_{0})-f(\theta^{*})}{\varepsilon}\right)$ in the definition of $\widetilde{T}$ corresponds to the number of outer iterations needed to obtain a sufficiently tight lower bound for the targeted accuracy. Similar order guarantees are established in Theorems 1 and 2. The proof of Theorem 2 is provided in Appendix B.
@@ 6 Numerical experiments
(i) Comparison to fixed step-size: In Figure 2 we present a comparison between IHT with a fixed step size (blue) and the adaptive step size used in Algorithm 1 (red) when solving linear regression (left panel) and logistic regression (center panel) respectively. When working with a constant step-size, we set the step size to $\frac{2}{3\bar{L}}$ following Jain et al. (2014) for both linear and logistic regression. $\bar{L}$ for linear regression can be upper bounded as $\lambda_{\max}(\Sigma)(3+\frac{2(2s+s^{*})}{s\alpha})$ (Loh and Wainwright, 2015, Appendix D.1), whereas $\bar{L}$ for logistic regression is one fourth of that of linear regression. In both settings, $\lambda_{\max}(\Sigma)\leq\frac{2}{(1-\omega)^{2}(1+\omega)}$ Agarwal et al. (2012). Although both step size strategies share the same theoretical guarantees, we observe that the adaptive step size speeds-up convergence. This advantage arises because an adaptive step size can adapt to the local curvature and therefore be significantly more aggressive, allowing for faster progress towards the solution. As expected, this effect is more pronounced when solving logistic regression where the functions’s curvature will depend on the point.
@@
A Main Theorems .A
 A.1 Proof of Theorem 1.A.1
 A.2 Proof of Corollary 1.A.2
 A.3 Proof of Corollary 2 .A.3
 A.4 Proof of Corollary 3 .A.4
B Adaptive Lower Bound .B
 B.1 Proof of Theorem 2.B.1
C Other Statistical Guarantees .C
 C.1 Sparse Linear Regression.C.1
 C.2 Generalized Linear Models.C.2
D Experiments on real data .D
 D.1 Linear Regression.D
 D.2 Logistic Regression.D
@@ A.2 Proof of Corollary 1
and for any $i\in\widehat{\mathcal{S}}$ we have
$|[\theta_{t}]_{i}-\gamma_{t}[g_{t}]_{i}|>\frac{\|\mathrm{HT}_{s}(\widehat{g})\|}{2\bar{\mu}}.$

### v2

@@ Abstract
We propose and study Sparse Polyak, a variant of Polyak’s adaptive step size, designed to solve high-dimensional statistical estimation problems where the problem dimension is allowed to grow much faster than the sample size. In such settings, the standard Polyak step size performs poorly, requiring an increasing number of iterations to achieve optimal statistical precision-even when, the problem remains well conditioned and/or the achievable precision itself does not degrade with problem size. We trace this limitation to a mismatch in how smoothness is measured: in high dimensions, it is no longer effective to estimate the Lipschitz smoothness constant. Instead, it is more appropriate to estimate the smoothness restricted to specific directions relevant to the problem (restricted Lipschitz smoothness constant). Sparse Polyak overcomes this issue by modifying the step size to estimate the restricted Lipschitz smoothness constant. We support our approach with both theoretical analysis and numerical experiments, demonstrating its improved performance.
@@ 3 Main Result
This theorem focuses on the distance of the iterates to $\theta^{*}$. The quantity $T$ can be interpreted as the number of iterations required to reach the desired accuracy when applying Algorithm 1 with $\widehat{f}=f^{*}$. The additional term $\mathcal{O}\left(\frac{f(\theta_{0})-f(\theta^{*})}{\varepsilon}\right)$ in the definition of $\widetilde{T}$ corresponds to the number of outer iterations needed to obtain a sufficiently tight lower bound for the targeted accuracy. Similar order guarantees are established in Theorems 1 and 2. The proof of Theorem 2 is provided in Appendix B.
@@ 6 Numerical experiments
(i) Comparison to fixed step-size: In Figure 2 we present a comparison between IHT with a fixed step size (blue) and the adaptive step size used in Algorithm 1 (red) when solving linear regression (left panel) and logistic regression (center panel) respectively. When working with a constant step-size, we set the step size to $\frac{2}{3\bar{L}}$ following Jain et al. (2014) for both linear and logistic regression. $\bar{L}$ for linear regression can be upper bounded as $\lambda_{\max}(\Sigma)(3+\frac{2(2s+s^{*})}{s\alpha})$ (Loh and Wainwright, 2015, Appendix D.1), whereas $\bar{L}$ for logistic regression is one fourth of that of linear regression. In both settings, $\lambda_{\max}(\Sigma)\leq\frac{2}{(1-\omega)^{2}(1+\omega)}$ Agarwal et al. (2012). Although both step size strategies share the same theoretical guarantees, we observe that the adaptive step size speeds-up convergence. This advantage arises because an adaptive step size can adapt to the local curvature and therefore be significantly more aggressive, allowing for faster progress towards the solution. As expected, this effect is more pronounced when solving logistic regression where the functions’ curvature will depend on the point.
@@
A  Main Theorems ........................................................................................................................................................................A
 A.1 Proof of Theorem 1........................................................................................................................................................................A.1
 A.2 Proof of Corollary 1........................................................................................................................................................................A.2
 A.3 Proof of Corollary 2 ........................................................................................................................................................................A.3
 A.4 Proof of Corollary 3 ........................................................................................................................................................................A.4
B  Adaptive Lower Bound ........................................................................................................................................................................B
 B.1 Proof of Theorem 2........................................................................................................................................................................B.1
C  Other Statistical Guarantees ........................................................................................................................................................................C
 C.1 Sparse Linear Regression........................................................................................................................................................................C.1
 C.2 Generalized Linear Models........................................................................................................................................................................C.2
D  Experiments on real data ........................................................................................................................................................................D
 D.1 Linear Regression........................................................................................................................................................................D
 D.2 Logistic Regression........................................................................................................................................................................D
@@ A.2 Proof of Corollary 1
and for any $i\in\widehat{\mathcal{S}}$, given that $\|\theta_{t}-\widehat{\theta}\|\leq\frac{6\|\mathrm{HT}_{s}(\widehat{g})\|}{\bar{\mu}}$, we have
$|[\theta_{t}]_{i}-\gamma_{t}[g_{t}]_{i}|>\frac{7\|\mathrm{HT}_{s}(\widehat{g})\|}{\bar{\mu}}-\frac{6\|\mathrm{HT}_{s}(\widehat{g})\|}{2\bar{\mu}}-\frac{\|\mathrm{HT}_{s}(\widehat{g})\|}{2\bar{\mu}}=\frac{\|\mathrm{HT}_{s}(\widehat{g})\|}{2\bar{\mu}}.$